OpenStudy (anonymous):

Use the substitution method to solve the system of equations. x + 2y = 15 and 2x - 3y = 2

OpenStudy (anonymous):

subtract 2y on both sides x=-2y+15 and plug in on the second equation for x so will be 2(-2y+15)-3y=2

OpenStudy (anonymous):

and solve for y

OpenStudy (anonymous):

-2(x+2y=15) 2x-3y=2 --------------- -2x-4y=-30 2x-3y=2 (add them) ------------------ -7y=-28 (divide by -7) ------------- y= 4 then substitute y in one of the equations --------------------------------- 2x-3(4)=2 2x-7=2 (add 7) 2x=9 x= 9/2 (then check)